A bag full of mangoes was divided into two heaps in the ratio 11:13. If the first heap has 143 mangoes what is the number of mangoes in the second heap?
Answer by ptaylor(2198)

Let x=total number of mangoes
If the ratio was 11:13, then the first heap has 11/(11+13)=11/24 and the other heap had 13/24 of the mangoes.
Soooo
(11/24)*x=143
x=143*24/11=312----total number of mangoes
312-143=169----------number of mangoes in second heap
or(13/24)*312=169
Hope this helps---ptaylor